Problem 4. = (1 point) Find the solution to the linear system of differential equations { -17x + 14 y -21x + 18y satisfying t(1 point) Find the most general real-valued solution to the linear system of differential equations = [-* ____* x1 (1) = C1 +STRUGGLING WITH THESE TWO SETS OF PROBLEMS ID APPREACIATE THE CORRECT ANSWERS THANKS
Answer #1

boly o conside A = 1-17 14 21 18 & IA-I ) = 1² -12= ((143) (1-4) ا وا ر3 = A د consider I A-1, I1=1-14 -) 14 21 RI SRL -14 1,